OpenStudy (anonymous):

you earn$7.25 per hour baby-sitting write and solve an inequality to find how many full hours you must work to earn atleast $200

OpenStudy (anonymous):

with like how you did it please

OpenStudy (mertsj):

7.25h >- 200

OpenStudy (anonymous):

27.5

OpenStudy (agreene):

7.25 per hour or: 7.25/h We want how many hours for AT LEAST 200 so: \[\frac{7.25}{h} \ge 200\] now it asks to solve it... 7.25/h >= 200 7.25 >= 200h 7.25/200 >= h 29/800 >= h Or in decimal: \[h \le 0.03625\]

OpenStudy (mertsj):

That is totally ridiculous, agreene. Think about it. If you work less than or equal to .036 hours, you will earn $200?

OpenStudy (agreene):

Yeah, I didnt really think that through, eh?

OpenStudy (anonymous):

@mertsj your right but how do you get that answer

OpenStudy (agreene):

I'd like to point out this question is worded incorrectly... and is generally stupid. 7.25h >= 200 h>= 200/7.25 h>= 27.586

OpenStudy (anonymous):

I think i have already gave her the answer...

OpenStudy (anonymous):

i copied word from word and doubled checked

OpenStudy (anonymous):

u got it wrong tho i need bio help

OpenStudy (mertsj):

I thought : if I work for $5 an hour in 2 hours I would make 10 dollars. In 3 hours I would make 15 dollars in 4 hours I would make 20 dollars. Then it dawned on me that if I want to know how much I make, I should multiply the hours I work by how much I make in an hour. In your problem, you didn't know how many hours you were working so I used the variable h for that. That's how I got 7.50h. Then the problem said you were going to make at least $200 which means you might make $200 or you might make more than $200. So: you make 7.50h in h hours you make more than or equal to 200 7.50h >-$200

OpenStudy (agreene):

I wasn't blaming you sugarpie, I just meant whoever wrote the question is obviously unaware of mathematical nomenclature.

OpenStudy (agreene):

When you say 7.25 per hour, that implies 7.25$/h or 7.25/h And not that you worked 7.25 hours which is what they really meant.

OpenStudy (anonymous):

ok thank you

OpenStudy (mertsj):

sugarpie, don't pay any attention to agreene. The problem is a perfectly fine problem. He just doesn't understand it either.

OpenStudy (anonymous):

alright lolcan you also help me with working with sets (i know u bet its easy but im in seventh grade taking a 9th grade course and i miissed that class and i have a test tomarrow)

OpenStudy (agreene):

@Mertsj I understand the problem. I just didn't really think it through in the way you did... I thought of it how it was written... once you pointed out my result made no sense in relation to the question asked... I looked what was wrong, and it was the question asking me to use quotients when it meant to state something rather different.

OpenStudy (anonymous):

OK lets not overreact here just please someone help me with sets